A 3 kg block is placed on top of a 7 kg block. The coefficient of kinetic friction between the 6 kg block and the surface is 0.3. A rightward force F is applied to the 6. kg block. Calculate the magnitude of the force necessary to pull both blocks to the right with an acceleration of 3 m/s^2.
1
Expert's answer
2020-11-16T07:53:16-0500
"(m+M)a=F-\mu N\\(m+M)a=F-\mu (m+M)g\\F=(m+M)(a+\mu g)\\F=(3+7)(3+0.3(9.8))=59.4\ N"